OpenStudy (anonymous):

Simplify. Write your answers without exponents. (1/64) ^ -2/3 25 ^ -3/2

OpenStudy (ybarrap):

(1/64)^(-2/3) = 64^(2/3) = (2^6)^(2/3)=2^(6*2/3)=2^(4)=16 25^(-3/2)=(1/25)^(3/2)=(1/5^2)^(3/2)=1/5^(2*3/2)=1/5^3= 1/125 Make sense?

OpenStudy (anonymous):

Ooh, Okay. It makes more sense then what I had in mind. Thank you! (: